A coil inductance 8.44 mH and resistance 6 Ω is connected to a. 12 V battery. The current in the coil is 1.0  A at approximately the time
(a) 500 s
(b) 25 s
(c) 35 ms
(d) 1 ms

Grade:11

1 Answers

Kevin Nash
9 years ago
Hello Student,
Please find the answer to your question
. (d) KEY CONCEPT : Using I = I0
But I0 = V/R and t = L/R
∴ I = V/R (1 – e-Rt/L) = 12/6 [ 1 – e-6t/8.4x 10-3]
1 (given)
∴ t = 0.97 x 10-3≈ 1 ms
(D)
1. (a,b,c)
1/RC, R/L and 1/ √LC have the dimensions of frequency.
Thanks
Kevin Nash
askIITians Faculty